OpenStudy (anonymous):

Use Cramer's rule to solve the system of equations. Write your answer as an ordered pair. Example: (2, -4) ... x-2y= -9 3x+y=1

OpenStudy (anonymous):

find determinant first....

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

lol "yea" is not a determinant..i said did u calculate the determinant. you know how to do it?

OpenStudy (anonymous):

ok fine see i will teach you. cramer rule says x = Dx / D y = Dy / D first you want to find D  so your D=1*1 - 2*3 = 1 - 6 = -5